OpenStudy (calculusxy):

Sally spent half of her allowance going to the movies. She washed the family car and earned 9 dollars. What is her weekly allowance if she ended with 18 dollars? Write an equation, then evaluate it. @Lena772 @Mel98 @Loser66 @agent0smith

OpenStudy (lena772):

Let x represent her allowance

OpenStudy (lena772):

x/2 is what she has left after the movies, agree?

OpenStudy (lena772):

Is the $9 an extra amount of money she earned or a part of her typical allowance?

OpenStudy (calculusxy):

I believe what she eanrned that week.

OpenStudy (lena772):

\[\frac{ x }{ 2 }+9=18\]

OpenStudy (lena772):

agree?

OpenStudy (calculusxy):

Ok go on..

OpenStudy (lena772):

\[\frac{ x }{ 2 }+9-9=18-9\]

OpenStudy (lena772):

\[\frac{ x }{ 2}=9\]

OpenStudy (calculusxy):

Would that be 4.8?

OpenStudy (lena772):

\[\frac{ x }{ 2 }*2=9*2\]

OpenStudy (lena772):

\[x=18\]

hero (hero):

Sorry to interrupt but @Lena772, when you get a chance, please re-visit the question about sprinkling a lawn

OpenStudy (lena772):

Okay, I will do than now @Hero

OpenStudy (lena772):

Do you understand @calculusxy

OpenStudy (calculusxy):

Yes thank you so much!

OpenStudy (lena772):

You're welcome if you ever need help just tag me

OpenStudy (lena772):

@Hero tag me in the sprinkler I can't find it in my questions :(